In August 2017, Richard purchased and placed in service an office building costing $753000 including $134000 for the land. The amount of depreciation Richard May claim in 2017 is

A) 5961
B) 7286
C) 7366
D) 8443




□Mark for follow up Question 30 of 75 Which of the following distributions from a Coverdell Education Savings Account (ESA) is most likely to be taxa O A distribution that is a return of investment, used to pay qualifying education O A distribution taken to pay education expenses already reduced by the American O A distribution used to pay tuition at a private high school. O A distribution that exceeds qualifying education expenses. expenses at an eligible institution Mark for follow up Back Save/Return Later Summary Next
Answer #1

1.Ans: A) $ 5961

Solution: Amount of Building = $ 753000-134000= $ 619000

Depreciable Period = 39 years

Depreciation Value per Year= $ 619000/39 years   =$15871.79

Using half year convention depreciation for 2016 = $ 15871.79x4.5/12   = $ 5961

Note:

1.As per MACRS, The useful life of Non- Residential real property is 39 Years

2. Depreciation is straight line method Using mid Month Convention which means, The month of acquisition is used as half month.

2)
Ans:
D)A distribution that exceeds qualifying education expenses

Note:
As Per the Provision of IRS,Distribution in excess of qualified education expenses or distribution not used for qualified education expenses will cause the earnings to be included in ordinary income.